OpenStudy (x3_drummerchick):

will give medals! - if you an explain how to solve this: find all fourth roots of -256i (this is my example problem)

OpenStudy (x3_drummerchick):

@zepdrix

OpenStudy (x3_drummerchick):

@mathmale

zepdrix (zepdrix):

So umm, are you familiar with Euler's Identity? You've seen complex numbers in both trig and exponential form, yes?\[\large\rm e^{i \theta}=\cos \theta+i \sin \theta\]

OpenStudy (x3_drummerchick):

no, but i was given demoires theorum: if z=r( (cosθ + i sinθ), then z^n= r^n( cos n θ + i sin n θ)

zepdrix (zepdrix):

No exponential? :) Aw ok fine fine.

OpenStudy (x3_drummerchick):

i havent learned the one you've posted

OpenStudy (x3_drummerchick):

sorry :( lol

zepdrix (zepdrix):

\[\large\rm z=-256i\]Taking our 4th root gives us,\[\large\rm z^{1/4}=\left(-256i\right)^{1/4}\]\[\large\rm z^{1/4}=256^{1/4}\left(-i\right)^{1/4}\]\[\large\rm z^{1/4}=4(-i)^{1/4}\]Let's take it one step further, and write our complex number like this:\[\large\rm z^{1/4}=4(0-i)^{1/4}\]Check those steps out a sec :) Lemme know if confusion sets in and rots your brain.

OpenStudy (x3_drummerchick):

okay that makes sense

zepdrix (zepdrix):

What we have in the brackets is some complex number which we can write in trig form. \[\large\rm z^{1/4}=4(~~\color{orangered}{0}~+\color{royalblue}{-i}~~)^{1/4}\]\[\large\rm z^{1/4}=4(~~\color{orangered}{\cos \theta}~+\color{royalblue}{i \sin \theta}~~)^{1/4}\] Can you figure out which angle... when you take the cosine ... gives you zero sine ... gives you -1

OpenStudy (x3_drummerchick):

no no, this makes sense, youre showing me how the cosine and sine are derived

zepdrix (zepdrix):

So I guess our angle is going to be 3pi/2 or -pi/2. Can you visualize that?